### v1

@@ 1 Introduction
Recent advances in time series foundation models (TSFMs) [goswami2024moment, rasul2023lag, liutimer2024timer, woo2024unified, liu2025sundial, arango2025chronosx], built upon large and diverse temporal corpora, have greatly advanced generalizable and adaptive modeling across domains such as finance, healthcare, and climate science. However, the performance of these large models fundamentally depends on the quality of their training data [wangoptimal, fang2023bayotide], making time series data valuation an indispensable component of TSFM development. Time series data valuation aims to quantify the contribution of individual time series samples to the overall model performance, enabling principled data selection, quality-aware training, and efficient resource allocation [sun20242d, nguyen2023context, zhang2025fairshare]. Accurate data valuation helps identify informative and trustworthy time series samples while filtering out corrupted or unrepresentative ones, ultimately improving both data efficiency and model performance [kwon2023data]. As TSFMs continue to expand in scale and scope, it becomes a compelling task to develop effective and scalable data valuation methods tailored to TSFMs.
Existing Time series data valuation methods, including TimeInf [zhang2024timeinf] and TimeShap [bento2021timeshap], extend theoretically grounded influence functions [hampel1974influence] and Shapley-values [shapley1953value] to temporal settings. These methods can effectively capture the temporal dependence nature of time series data [kunsch1984infinitesimal] by incorporating temporal segmentation and sequence-aware analysis. Despite their effectiveness and theoretical foundations, these methods unfortunately inherit the computational burdens of their underlying theory: influence function-based approaches require costly Hessian and gradient calculations, while Shapley-based methods incur exponential subset sampling. Consequently, although effective for small-scale linear or recurrent models, these approaches are infeasible for large-scale TSFMs that typically come with millions or billions of parameters.
@@ 2.1 Time Series Foundation Models
Recent advancements in time series foundation models (TSFMs) have significantly reshaped the landscape of time series analysis and modeling. Similar to large language models (LLMs), TSFMs are pretrained on massive and heterogeneous time series corpora to learn universal temporal representations that can be adapted to various downstream tasks. These models adopt diverse architectures, including encoder-only Transformers [goswami2024moment, woo2024unified], decoder-only autoregressive Transformers [garza2023timegpt, liutimer2024timer, das2024decoder, liu2024timer], and mixture-of-experts based sparse structures [xiaoming2025time, liu2024moirai, wu2025unlocking]. Model sizes now range from tens of millions to over 2.4 billion parameters [xiaoming2025time], trained on tens of billions of temporal observations collected from diverse real-world domains.
While this scaling has led to increasingly powerful and versatile temporal representations, it also amplifies the dependency of TSFM performance on high-quality time series data. Low-quality or biased time series samples can degrade the learned representations and reduce downstream generalization ability [zhang2024irregular, liu2024time]. Therefore, accurately valuing and selecting high-quality time series data for TSFMs becomes a fundamental yet underexplored problem, motivating the development of our proposed LTSV.
@@ 2.2 Data Valuation Methods
Time Series Data Valuation. Time series data valuation requires specialized techniques that explicitly consider temporal dependencies and correlations. Building upon classical principles such as the influence function [hampel1974influence] and Shapley value [shapley1953value], several methods have been proposed to extend data attribution into the time series domain. For example, TimeInf [zhang2024timeinf] adapts the influence function to temporal settings by modeling the effect of perturbing autoregressive processes, while TimeShap [bento2021timeshap] and WindowShap [nayebi2023windowshap] generalize Shapley-based attribution through sliding window decomposition and temporal feature aggregation. Subsequent studies extend influence estimation to multivariate settings [wang2024channel] and incorporate sequence-aware Shapley computations into Transformer architectures [cheng2025unifying]. Despite these advances, both influence-function and Shapley-based approaches remain computationally expensive, requiring either costly Hessian inversions or exponential subset sampling, and thus making them impractical for large-scale TSFMs with millions of parameters.
Data Valuation for Foundation Models. Beyond time series–specific approaches, data valuation on large foundation models has been widely explored in the language domain. Existing works employ strategies such as rule-based scoring [penedo2024fineweb], semantic deduplication metrics [penedo2023refinedweb, tirumala2023d4], and proximity-based quality assessment via perplexity [du2022glam] or embedding similarity [xie2023data], while recent advances further leverage powerful reference LLMs for quality estimation [wettig2024qurating, sachdeva2024train, yu2024mates]. However, these methods rely on text-centric characteristics, making them ill-suited for TSFMs characterized by non-stationary, correlated, and multi-channel dynamics. To address these gaps, our LTSV framework leverages the in-context fine-tuning to approximate influence-based valuation in a lightweight and temporally aware manner.
@@ 2.3 Preliminaries
We revisit the classical Influence Function [koh2017understanding], which serves as the theoretical basis for our proposed time series data valuation framework.
@@
The derivation follows the first-order influence approximation in [pruthi2020estimating, koh2017understanding].
This classical formulation provides a principled mechanism to quantify how a single target sample affects the model’s prediction or context loss. However, computing the inverse Hessian $H_{\theta^{*}}^{-1}$ is computationally infeasible for large-scale deep or foundation models [basu2020influence], which motivates the development of more efficient data valuation approaches such as our proposed LTSV.
@@ 3.1 Influence Function Approximation via In-Context Finetuning
A single-step gradient update on the target sample $z$ gives $\theta_{\mathrm{finetuned}}=\theta-\eta\nabla_{\theta}\mathcal{L}(x,y;\theta)$, where $\eta$ is the learning rate. This first-order approximation is valid for a sufficiently small $\eta$ [pruthi2020estimating]. Substituting $\Delta\theta=\theta_{\mathrm{finetuned}}-\theta\approx-\eta\nabla_{\theta}\mathcal{L}(x,y;\theta)$ into Eq.(1) yields
@@
Therefore, the influence of a target sample can be locally approximated by the loss reduction on context samples after one-step in-context finetuning [iter2023context, yu2024mates]. This equivalence provides theoretical evidence for approximating the computationally demanding influence function with more computationally efficient approximations via in-context finetuning and context loss calculation, forming the foundation of our lightweight time series data valuation methods for TSFMs presented below.
@@ 3.2 Time Series Data Valuation on TSFM via In-Context Finetuning
Block-Level Time Series Data Value Estimation. Equipped with theoretical support for the approximation capability of in-context finetuning to influence functions, we describe the detailed algorithm for tailoring in-context finetuning to time series data valuation. Specifically, we further divide the target time series dataset $\mathcal{D}$ into overlapping blocks $\{B_{k}\}_{k=1}^{K}$ of fixed length $L$, preserving local temporal dependencies [hall1985resampling, buhlmann2002bootstraps].
@@
Each block $B_{k}$ is treated as a data valuation unit. We apply in-context finetuning to TSFMs (e.g., Time-MoE [xiaoming2025time], a 2.4B-parameter model) by treating $B_{k}$ as the target sample and $\mathcal{D}_{\text{context}}$ as the context samples, thereby approximating the influence function value of $B_{k}$ by calculating the context loss following Eq.(3):
@@ 4.1 Experimental Setup
Datasets, TSFM Types and Baselines. We evaluate our proposed LTSV on five widely used time series forecasting benchmarks that cover diverse temporal patterns and domains: Electricity [trindade2015electricityloaddiagrams20112014], Exchange Rate [lai2018modeling], Weather [weather_data], ETT [zhou2021informer](ETT-m2), and Illness [illness_data]. Experiments are conducted on three representative time series foundation models that cover different architectural paradigms of TSFMs: Time-MoE [woo2024unified], a mixture-of-experts decoding-only Transformer model; Time-LLM [liutimer2024timer], which integrates large language model architectures; and MOMENT [goswami2024moment], a T5-based encoder-decoder Transformer model. We compare LTSV against representative data valuation methods, including Influence Function [zhang2024timeinf], and Data Shapley [bento2021timeshap].
@@ 4.4 Generalization of Foundation Model Valuations to Diverse Downstream Models
In this section, we investigate whether the data valuations obtained from the foundation model can generalize effectively to diverse downstream time series models. Specifically, we train each downstream model using five subsets of data: (1) the top-valued samples according to LTSV scores, (2) the bottom-valued samples from LTSV, (3) the samples ranked highest by the classical influence function adapted for time series [zhang2024timeinf], (4) the samples ranked highest by the Shapley value method (time series version) [bento2021timeshap], and (5) randomly selected samples. The selection ratio is fixed at 20% to better highlight the performance differences among methods. The data valuations are obtained from the Time-MoE foundation model. The downstream forecasting models cover a wide range of architectures, including DLinear [zeng2023transformers], PatchTST [nie2022time], and PAttn [tan2024language], representing linear, transformer-based, and attention-based paradigms respectively.
@@
Influence [zhang2024timeinf] 2.074 1.310 0.360 0.522 0.322 0.478 1.718 1.142 0.355 0.496
Shapley [bento2021timeshap] 2.047 1.301 0.372 0.529 0.384 0.520 1.603 1.102 0.325 0.475
@@
Influence [zhang2024timeinf] 1.830 1.131 0.286 0.456 0.355 0.508 1.099 0.861 0.314 0.459
Shapley [bento2021timeshap] 1.888 1.145 0.288 0.460 0.366 0.516 1.057 0.844 0.266 0.411
@@
Influence [zhang2024timeinf] 1.883 1.139 0.308 0.462 0.326 0.487 1.025 0.841 0.326 0.468
Shapley [bento2021timeshap] 1.925 1.153 0.313 0.478 0.474 0.574 0.965 0.806 0.292 0.432
@@ 4.5 Ablation Study on Block Length
Influence function [zhang2024timeinf] 1.281 0.994 1.678 1.164 2.074 1.310 2.017 1.277

### v2

@@ Abstract
22footnotetext: This paper has been accepted as a full paper in DASFAA 2026.
@@ 1 Introduction
Recent advances in time series foundation models (TSFMs) [10, 28, 22, 39, 21, 1], built upon large and diverse temporal corpora, have greatly advanced generalizable and adaptive modeling across domains such as finance, healthcare, and climate science. However, the performance of these large models fundamentally depends on the quality of their training data [35, 8], making time series data valuation an indispensable component of TSFM development. Time series data valuation aims to quantify the contribution of individual time series samples to the overall model performance, enabling principled data selection, quality-aware training, and efficient resource allocation [31, 24, 45]. Accurate data valuation helps identify informative and trustworthy time series samples while filtering out corrupted or unrepresentative ones, ultimately improving both data efficiency and model performance [17]. As TSFMs continue to expand in scale and scope, it becomes a compelling task to develop effective and scalable data valuation methods tailored to TSFMs.
Existing Time series data valuation methods, including TimeInf [47] and TimeShap [3], extend theoretically grounded influence functions [12] and Shapley-values [30] to temporal settings. These methods can effectively capture the temporal dependence nature of time series data [16] by incorporating temporal segmentation and sequence-aware analysis. Despite their effectiveness and theoretical foundations, these methods unfortunately inherit the computational burdens of their underlying theory: influence function-based approaches require costly Hessian and gradient calculations, while Shapley-based methods incur exponential subset sampling. Consequently, although effective for small-scale linear or recurrent models, these approaches are infeasible for large-scale TSFMs that typically come with millions or billions of parameters.
@@ 2.1 Time Series Foundation Models
Recent advancements in time series foundation models (TSFMs) have significantly reshaped the landscape of time series analysis and modeling. Similar to large language models (LLMs), TSFMs are pretrained on massive and heterogeneous time series corpora to learn universal temporal representations that can be adapted to various downstream tasks. These models adopt diverse architectures, including encoder-only Transformers [10, 39], decoder-only autoregressive Transformers [9, 22, 6], and mixture-of-experts based sparse structures [41, 20, 40]. Model sizes now range from tens of millions to over 2.4 billion parameters [41], trained on tens of billions of temporal observations collected from diverse real-world domains.
While this scaling has led to increasingly powerful and versatile temporal representations, it also amplifies the dependency of TSFM performance on high-quality time series data. Low-quality or biased time series samples can degrade the learned representations and reduce downstream generalization ability [46, 19]. Therefore, accurately valuing and selecting high-quality time series data for TSFMs becomes a fundamental yet underexplored problem, motivating the development of our proposed LTSV.
@@ 2.2 Data Valuation Methods
Time Series Data Valuation. Time series data valuation requires specialized techniques that explicitly consider temporal dependencies and correlations. Building upon classical principles such as the influence function [12] and Shapley value [30], several methods have been proposed to extend data attribution into the time series domain. For example, TimeInf [47] adapts the influence function to temporal settings by modeling the effect of perturbing autoregressive processes, while TimeShap [3] and WindowShap [23] generalize Shapley-based attribution through sliding window decomposition and temporal feature aggregation. Subsequent studies extend influence estimation to multivariate settings [36] and incorporate sequence-aware Shapley computations into Transformer architectures [5]. Despite these advances, both influence-function and Shapley-based approaches remain computationally expensive, requiring either costly Hessian inversions or exponential subset sampling, and thus making them impractical for large-scale TSFMs with millions of parameters.
Data Valuation for Foundation Models. Beyond time series–specific approaches, data valuation on foundation models has been widely explored in the NLP domain. Existing works employ strategies such as rule-based scoring [26], semantic deduplication metrics [33], and proximity-based quality assessment via perplexity [7] or embedding similarity [42], while recent advances leverage powerful reference LLMs for quality estimation [38, 29, 43]. However, these methods rely on text-centric characteristics, making them ill-suited for TSFMs characterized by non-stationary, correlated, and multi-channel dynamics. To address these gaps, our LTSV framework leverages the in-context fine-tuning to approximate influence-based valuation in a lightweight and temporally aware manner.
@@ 2.3 Preliminaries
We revisit the classical Influence Function [15], which serves as the theoretical basis for our proposed time series data valuation framework.
@@
The derivation follows the first-order influence approximation in [27, 15].
This classical formulation provides a principled mechanism to quantify how a single target sample affects the model’s prediction or context loss. However, computing the inverse Hessian $H_{\theta^{*}}^{-1}$ is computationally infeasible for large-scale deep or foundation models [2], which motivates the development of more efficient data valuation approaches such as our proposed LTSV.
@@ 3.1 Influence Function Approximation via In-Context Finetuning
A single-step gradient update on the target sample $z$ gives $\theta_{\mathrm{finetuned}}=\theta-\eta\nabla_{\theta}\mathcal{L}(x,y;\theta)$, where $\eta$ is the learning rate. This first-order approximation is valid for a sufficiently small $\eta$ [27]. Substituting $\Delta\theta=\theta_{\mathrm{finetuned}}-\theta\approx-\eta\nabla_{\theta}\mathcal{L}(x,y;\theta)$ into Eq.(1) yields
@@
Therefore, the influence of a target sample can be locally approximated by the loss reduction on context samples after one-step in-context finetuning [14, 43]. This equivalence provides theoretical evidence for approximating the computationally demanding influence function with more computationally efficient approximations via in-context finetuning and context loss calculation, forming the foundation of our lightweight time series data valuation methods for TSFMs presented below.
@@ 3.2 Time Series Data Valuation on TSFM via In-Context Finetuning
Block-Level Time Series Data Value Estimation. Equipped with theoretical support for the approximation capability of in-context finetuning to influence functions, we describe the detailed algorithm for tailoring in-context finetuning to time series data valuation. Specifically, we further divide the target time series dataset $\mathcal{D}$ into overlapping blocks $\{B_{k}\}_{k=1}^{K}$ of fixed length $L$, preserving local temporal dependencies [11, 4].
@@
Each block $B_{k}$ is treated as a data valuation unit. We apply in-context finetuning to TSFMs (e.g., Time-MoE [41], a 2.4B-parameter model) by treating $B_{k}$ as the target sample and $\mathcal{D}_{\text{context}}$ as the context samples, thereby approximating the influence function value of $B_{k}$ by calculating the context loss following Eq.(3):
@@ 4.1 Experimental Setup
Datasets, TSFM Types and Baselines. We evaluate our proposed LTSV on five widely used time series forecasting benchmarks that cover diverse temporal patterns and domains: Electricity [34], Exchange Rate [18], Weather [37], ETT [48](ETT-m2), and Illness [13]. Experiments are conducted on three representative time series foundation models that cover different architectural paradigms of TSFMs: Time-MoE [39], a mixture-of-experts decoding-only Transformer model; Time-LLM [22], which integrates large language model architectures; and MOMENT [10], a T5-based encoder-decoder Transformer model. We compare LTSV against representative data valuation methods, including Influence Function [47], and Data Shapley [3].
@@ 4.4 Generalization of Foundation Model Valuations to Diverse Downstream Models
In this section, we investigate whether the data valuations obtained from the foundation model can generalize effectively to diverse downstream time series models. Specifically, we train each downstream model using five subsets of data: (1) the top-valued samples according to LTSV scores, (2) the bottom-valued samples from LTSV, (3) the samples ranked highest by the classical influence function adapted for time series [47], (4) the samples ranked highest by the Shapley value method (time series version) [3], and (5) randomly selected samples. The selection ratio is fixed at 20% to better highlight the performance differences among methods. The data valuations are obtained from the Time-MoE foundation model. The downstream forecasting models cover a wide range of architectures, including DLinear [44], PatchTST [25], and PAttn [32], representing linear, transformer-based, and attention-based paradigms respectively.
@@
Influence [47] 2.074 1.310 0.360 0.522 0.322 0.478 1.718 1.142 0.355 0.496
Shapley [3] 2.047 1.301 0.372 0.529 0.384 0.520 1.603 1.102 0.325 0.475
@@
Influence [47] 1.830 1.131 0.286 0.456 0.355 0.508 1.099 0.861 0.314 0.459
Shapley [3] 1.888 1.145 0.288 0.460 0.366 0.516 1.057 0.844 0.266 0.411
@@
Influence [47] 1.883 1.139 0.308 0.462 0.326 0.487 1.025 0.841 0.326 0.468
Shapley [3] 1.925 1.153 0.313 0.478 0.474 0.574 0.965 0.806 0.292 0.432
@@ 4.5 Ablation Study on Block Length
Influence function [47] 1.281 0.994 1.678 1.164 2.074 1.310 2.017 1.277
@@ 5 Conclusion
{credits}
5.0.1 Acknowledgements
We would like to thank the three anonymous reviewers, the area chairs, and the program chairs for their constructive comments and efforts.